The length of a rectangle is 12 in. More than its width. The perimeter is 64 in. What is the length of the rectangle?
Answer by stanbon(75887)

The length of a rectangle is 12 in. More than its width. The perimeter is 64 in. What is the length of the rectangle?
-------------------
Equations:
l = w + 12
P = 2(length + width)
-----
64 = 2(w+12 + w)
32 = 2w+12)
2w = 20
width = 10 in
length = w + 12 = 22 in
-----------------------------
Cheers,
Stan H.
---------------